Hamburgular stole two fifths of ronald McDonald hamburgers .of the remaining hamburgers ,five sixths were still warm.if Ronald McDonald had 7 cold hamburgers left how many did Hamburgular steal?
Answer by rfer(16322)

7/(1/6)=
7*6/1=42 left
42/(3/5)=
42*5/3=210/3=70 to start
70*2/5=140/5=28 stolen